Fermi National Laboratory
 

Operation Summary for November 3, 2010 - November 5, 2010

 

Wednesday Nov. 3

The day shift began with the Tevatron (TeV) off, with Operations monitoring the Recycler stash (341.2E10), with the Antiproton Source (Pbar) stacking (35.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and D-Zero (D0) experiments in access, and with the Collider Detector at Fermilab (CDF) waiting for the next store.

During the day shift, D0 came out of access at 8:39 AM and Operations began the TeV low beta magnet turn on. At 10:38 AM, cryo system personnel reported that the TeV sector C1 wet engine needed a flywheel change. However, at 11:54 AM, they determined that the engine really needed a complete overhaul, which would take eight hours. Operators prepared the D0 and CDF collision halls for accesses as well as the TeV A-E sectors. Mechanical technicians repaired a LCW leak on a Pbar line magnet.

During the evening shift, the TeV accesses finished at 5 PM and Operations began the TeV check out. Unfortunately, the C1 cryo system quenched when the wet engine was brought online. Operators replaced a TeV corrector power supply that had tripped off. The Pbar Debuncher injection septa power supply tripped off on an over-temperature indication. Operations called in EE Support personnel to assist.

Thursday Nov. 4

The midnight shift began with operators checking the TeV orbits, with Operations monitoring the stash (490.2E10) and the Pbar stack (95.7E10), with Pbar off due to a Debuncher septa problem, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TeV checkout and Pbar repairs continued. During the TeV checkout, operators found and replaced a bad feed down sextupole magnet power supply. Operations put the TeV into HEP shot setup at 4:03 AM and established store 8242 at 5:10 AM with an average initial luminosity (AIL) of 322.9E30. EE Support personnel completed the Pbar septa repairs at 3:31 AM; they replaced a bad high voltage cable.

During the day shift, Operations terminated store 8242 at 10:57 AM, put the TeV into shot setup, and established store 8243 at 11:59 AM with an AIL of 345.5E30. Operators prepared Booster, Linac, NuMI Absorber, and MI for accesses that began around 12:28 PM. Vacuum technicians repaired a vacuum leak in Linac. Booster came out of access around 4 PM.

During the evening shift, Linac, NuMI Absorber, and MI were out of access by 6:04 PM. Operations established beam to MiniBooNE at 7:03 PM. A problem with the MI power supply loop controller delayed the MI beam. Power supply personnel came in and swapped out the loop controller; then a ground fault detector held off the MI turn on until the bypass card from the old controller was swapped with the card in the new controller. Operations established beam to Pbar and NuMI at 10:35 PM. Operations reported that several Booster RF trips occurred after beam returned.

Friday Nov. 5

The midnight shift began with Operations monitoring store 8243 (86.89E30) and the stash (35.4E10), with Pbar stacking (29.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to the returning Meson FTBF T-1004 experiment at 4:20 AM. Operators tuned the MI beam. Operations reported a couple of minor Linac RF system trips.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/08/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory